# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v2

@@ VI Future Research Directions
VI-A Energy-Efficient and Green LM Agents
With the increasingly widespread deployment of LM agents, their energy consumption and environmental impact have emerged as critical concerns. As reported, the energy consumed by ChatGPT to answer a single question for 590 million users is comparable to the monthly electricity usage of 175,000 Danes [90]. Given the exponential growth in model size and the computational resources required, energy-efficient strategies are essential for sustainable AI development, with the aim to reduce the significant carbon footprint associated with training and operating LM agents. Enabling technologies for energy-efficient and green LM agents include model compression techniques [26, 28], such as pruning, quantization, and knowledge distillation, which reduce the size and computational requirements of LMs without significantly affecting their accuracy. Additionally, the use of edge computing [70] and FL [85] allows for the distribution of computational tasks across multiple devices, thereby reducing the energy burden on central servers and enabling real-time processing with lower latency. Innovations in hardware [29], such as energy-efficient GPUs and TPUs, also play a critical role in achieving greener LM agents by optimizing the energy use of the underlying computational infrastructure.
@@ VI-A Energy-Efficient and Green LM Agents
VI-B Fair and Explainable LM Agents
As LM agents continue to play an increasingly central role in decision-making across various domains, the need for fairness and explainability becomes paramount to build trust among users, ensure compliance with ethical standards, and prevent unintended biases. It is particular for sensitive areas such as healthcare, finance, and law, where decisions should be transparent, justifiable, and free from bias. Bias detection and mitigation algorithms such as adversarial debiasing [208], reweighting [209], and fairness constraints [83] can be integrated into the training process to ensure that the models are less prone to propagate existing biases, thereby identifying and correcting biases in data and model outputs. Moreover, explainable AI (XAI) methods [210] such as shapley additive explanations (SHAP), local interpretable model-agnostic explanations (LIME), and counterfactual explanations allow users to understand the reasoning behind the model’s predictions, thereby enhancing trust, transparency, and accountability.
However, several key challenges remain be addressed. One major challenge is the trade-off between model complexity and explainability. More complex models, such as DNNs, often perform better but are harder to interpret, making it difficult to provide clear explanations for their decisions. Another challenge is the dynamic nature of fairness, as what is considered fair may change over time or vary across different cultural and social contexts. Ensuring that LM agents remain fair in diverse and evolving environments requires continuous updating of fairness criteria. Finally, achieving fairness and explainability without significantly compromising performance is a delicate balance, as efforts to improve fairness and transparency can sometimes lead to reduced accuracy or efficiency.
VI-C Cyber-Physical-Social Secure LM Agent Systems
@@
VI-D Value Ecosystem of LM Agents

### v3

@@ VII Future Research Directions
VII-A Energy-Efficient and Green LM Agents
With the increasingly widespread deployment of LM agents, their energy consumption and environmental impact have emerged as critical concerns. As reported, the energy consumed by ChatGPT to answer a single question for 590 million users is comparable to the monthly electricity usage of 175,000 Danes [95]. Given the exponential growth in model size and the computational resources required, energy-efficient strategies are essential for sustainable AI development, with the goal to lower the significant carbon footprint associated with training and operating LM agents. Enabling technologies for energy-efficient and green LM agents include model compression techniques [20, 22], such as quantization, pruning, and distillation, which lower model size and computational demands of LMs with minimal impact on accuracy. Additionally, the use of edge computing [66] and FL [87] enables the distribution of computational tasks across edge nodes, thereby reducing the energy burden on central servers and enabling real-time processing with lower latency. Innovations in hardware [23], such as energy-efficient GPUs and TPUs, also play a critical role in achieving greener LM agents by optimizing the energy use of the underlying computational infrastructure.
@@ VII-A Energy-Efficient and Green LM Agents
VII-B Fair and Explainable LM Agents
As LM agents continue to play an increasingly central role in decision-making across various domains, the need for fairness and explainability becomes paramount to build trust among users, ensure compliance with ethical standards, and prevent unintended biases. It is particular for sensitive areas such as healthcare, finance, and law, where decisions should be transparent, justifiable, and free from bias. Bias detection and mitigation algorithms, e.g., adversarial debiasing [210], reweighting [211], and fairness constraints [85], can be incorporated into the training process to ensure that the LM agents are less prone to propagate existing biases, thereby identifying and correcting biases in data and model outputs. Moreover, explainable AI (XAI) mechanisms [212], e.g., local interpretable model-agnostic explanations (LIME), shapley additive explanations (SHAP), and counterfactual explanations allow users to understand the reasoning behind the LM’s predictions, thereby enhancing trust, transparency, and accountability.
However, several key challenges remain to be addressed. One major challenge lies in balancing model complexity and explainability. While more complex models, such as DNNs, tend to achieve superior performance, they are often difficult to interpret. Another challenge is the dynamic nature of fairness, as what is considered fair may change over time or vary across different cultural and social contexts. Ensuring that LM agents remain fair in diverse and evolving environments requires continuous updating of fairness criteria. Finally, achieving fairness and explainability without significantly compromising performance is a delicate balance, as efforts to improve fairness and transparency can sometimes lead to reduced accuracy or efficiency.
VII-C Cyber-Physical-Social Secure LM Agent Systems
@@
VII-D Value Ecosystem of LM Agents